# Average Cost Optimality of Partially Observed MDPs: Contraction of Non-Linear Filters, Optimal Solutions and Approximations

### v2

@@ 2. Proof of the Main Theorem
    $\rho_{BL}\left(\eta(\cdot\mid z,u),\eta\left(\cdot\mid z^{\prime},u\right)\right)\leq 3\left(1+\alpha\right)\rho_{BL}\left(z,z^{\prime}\right).$
    for any $z,z^{\prime}\in{\mathcal{P}}(\mathbb{X})$.
@@
    for any $z,z^{\prime}\in{\mathcal{P}}(\mathbb{X})$.
@@
Under Assumption 1-4, we have that $\tilde{c}$ is Lipschitz continuous, since
@@
In view of Lemma 2.1, all conditions of Lemma 2.2 are satisfied. Assumption 9-2 follows from Theorem 2.3, and Assumption 9-5 follows from (13). Notably, the equi-continuity condition, Assumption 9-6 holds as a result of the relation
@@
for all $\beta\in(0,1)$. This concludes the proof.
@@ 3. Implications for Approximations, Robustness and Learning
In this section, we discuss the implications of our results, namely, that of the existence of a solution to the average cost optimality equation and Wasserstein regularity.
@@
We then study approximate optimality through the quantization of the belief space, or $P(\mathbb{X})$. This method yields a near-optimal policy for the original problem. We then present an alternate method to construct a finite Markov Decision Process (MDP). This is achieved by replacing the complete observable Markov process with the most recent N observations and actions. Within this finite state MDP framework, we apply Q-learning to obtain the optimal policy.
@@ 3.1. Robustness to Incorrect Priors for Discounted and Average Cost Criteria
However, they further refine these results by considering the effects of filter stability on robustness to incorrect priors.
@@
As mentioned in Note 3.2, [28], Theorem 4.1, demonstrates that if the condition $(1-\tilde{\delta}(T))(2-\delta(Q))<1$ is met, then the filter is universally stable in total variation in expectation.
@@
$J_{\beta}(\mu,\gamma^{\nu}_{\beta})-J^{*}_{\beta}(\mu)\leq\inf_{n\in\mathbb{N}}\left(\|c\|_{\infty}\frac{(1-\beta^{n})}{1-\beta}+\beta^{n}\frac{DK_{1}}{1-K_{2}\beta}+4\frac{\|c\|_{\infty}}{1-\beta}(\bar{\alpha}\beta)^{n}\right)$
Note that, unlike [26, Theorem 3.10], we are able to bound this expression even as $\beta\to 1$.
@@
where $\nu^{\prime}=\pi_{n-}^{\nu,\gamma_{\beta}^{\nu}}$ and $\gamma_{\beta}^{\nu^{\prime}}$ is an optimal discounted policy for $\nu^{\prime}$. As in [26], we can refer to these three components as transient cost, strategic measure cost, and approximation cost.
The transient cost (15) is upper bound by
@@ 3.2. Approximations and Learning
Throughout the rest of this section, we assume that Assumption 1 holds. Earlier, we have already shown that if Assumption 1 is valid for POMDP, then Assumption 9 is applicable for belief MDP. The following theorems assist in finding near-optimal policies for the average cost criteria; for further analysis see [29, Theorem 5] or [24, Theorem 7.3.6]. In particular, that a solution to the average cost optimality equation exists and this is arrived at via the vanishing discount method in our analysis earlier is critical for the applicability of the following result:
@@
Then, for every $\epsilon>0$, there exists sufficiently large $\beta$ such that
@@
This implies that the discounted cost optimal policy is near-optimal for the average cost criterion. b) Under Assumption 9, let $\beta_{\epsilon}$ be chosen as above, such that with $h_{\beta}$ as given in (14),
@@
Since Assumption 1 implies to Assumption 9, the above theorem also holds under Assumption 1. Therefore, under Assumption 1, a near-optimal policy for discounted cost criteria is also near-optimal for the average cost criteria. The subsequent subsections will detail how to establish a near-optimal policy for discounted costs.
@@ 3.2.1. Near Optimality of Finite Models via Quantization for Average Cost
  • During exploration, every possible state-action pair in $\bar{Z}\times\mathbb{U}$ is visited an infinite number of times [^1] This can be relaxed and refined to concern those states in the support of an invariant measure under an exploration policy..
@@ 3.2.2. Near Optimality of Finite Window Policies for Average Cost
$z_{n-N}^{-}$ $=\operatorname{Pr}\left(X_{n-N}\in\cdot\mid y_{n-N-1},\ldots,y_{0},u_{n-N-1},\ldots,u_{0}\right),$
$I_{n}^{N}$ =$=\left\{y_{n},\ldots,y_{n-N},u_{n-1},\ldots,u_{n-N}\right\},$
@@
With filter stability, as $N$ approaches infinity, the upper bound tends to zero. Given filter stability and a sufficiently large $N$, the policy $\tilde{\varphi}^{N}$ becomes near-optimal for the finite window belief MDP for discounted cost criteria. If $\beta$ is sufficiently large, this policy also becomes near-optimal for the average cost criterion.
@@
    During the exploration phase, every $(I,u)$ pair is visited infinitely often.
@@
[28], Theorem 4.1 provides sufficient conditions for uniform filter stability. If we have $\bar{\alpha}=(1-\tilde{\delta}(T))(2-\delta(Q))<1$, then the filter is exponentially stable with coefficient $\bar{\alpha}$ for any control policy. Here, $\tilde{\delta}(T)=\inf_{u\in\mathcal{U}}\delta(T(\cdot|\cdot,u))$.
@@ 4. Concluding Remarks
The average cost optimality is a challenging problem, especially under the average cost criterion. In this paper, we presented explicit and easily testable conditions for the existence of solutions to the average cost optimality equation where the state space is compact. A comparison with the related literature and several examples are presented. Notably, our paper appears to be the first one where a contraction result is presented and utilized for the optimality analysis. We finally presented several implications of our analysis and existence result on approximations, learning, and robustness.
@@
where $w_{y_{1}}=\left(z_{1}\left(z_{0}^{\prime},u,y_{1}\right)-z_{1}\left(z_{0},u,y_{1}\right)\right)$ which is a signed measure on $\mathbb{X}$. By the Measurable Selection Theorem choose measurable

### v3

@@ 2. Proof of the Main Theorem
    $\rho_{BL}\left(\eta(\cdot\mid z,u),\eta\left(\cdot\mid z^{\prime},u\right)\right)\leq 3\left(1+\alpha\right)\rho_{BL}\left(z,z^{\prime}\right)$
    for any $z,z^{\prime}\in{\mathcal{P}}(\mathbb{X})$ and $u\in\mathbb{U}$.
@@
    for any $z,z^{\prime}\in{\mathcal{P}}(\mathbb{X})$ and $u\in\mathbb{U}$, where $\tilde{\delta}(\mathcal{T}):=\inf_{u\in\mathbb{U}}\delta(\mathcal{T}(\cdot\mid\cdot,u))$
@@
Under Assumption 1-4, we have that $\tilde{c}$ defined in (3) is Lipschitz continuous, since
@@
First, we will show that if a partially observable Markov decision process (POMDP) satisfies Assumption 1, then the corresponding belief MDP satisfies Assumption 9.
Assumption 9-1 is valid due to equation (2). Assumptions 9-3 and 4 hold because $\mathbb{U}$ and $\mathcal{Z}$ are compact. Assumption 9-2 follows from Theorem 2.3, and Assumption 9-5 follows from (15).
By inequality (15), we know
@@
for all $\beta\in(0,1)$. Therefore, $h_{\beta}$ is equicontinuous for all $\beta\in(0,1)$. Since this condition holds for all subsequences, Assumption 9-6 is satisfied.
Thus, if a POMDP satisfies Assumption 1, then the belief MDP satisfies Assumption 9. All conditions of Lemma 2.2 are satisfied, and the proof is completed using Lemma 2.2.
@@ 3. Implications for Approximations, Robustness and Learning
In this section, we discuss several implications of our results on the existence of a solution to the average cost optimality equation and Wasserstein regularity.
@@
We then study approximate optimality through the quantization of the belief space, or $\mathcal{P}(\mathbb{X})$. This method yields a near-optimal policy for the original problem. We then present an alternate method to construct a finite Markov Decision Process (MDP). This is achieved by replacing the complete observable Markov process with the most recent N observations and actions. Within this finite state MDP framework, we apply Q-learning to obtain a near-optimal policy.
@@ 3.1. Robustness to Incorrect Priors for Discounted and Average Cost Criteria
However, by utilizing filter stability, the bound can be refined, as we will observe in the following.
@@
As mentioned in Note 3.2, [28], Theorem 4.1, demonstrates that if the condition $(1-\tilde{\delta}(\mathcal{T}))(2-\delta(Q))<1$ is met, then the filter is universally stable in total variation in expectation.
@@
$J_{\beta}(\mu,\gamma^{\nu}_{\beta})-J^{*}_{\beta}(\mu)\leq\inf_{n\in\mathbb{N}}\left(\|c\|_{\infty}\frac{(1-\beta^{n})}{1-\beta}+\beta^{n}\frac{DK_{1}}{1-K_{2}\beta}+4\frac{\|c\|_{\infty}}{1-\beta}(\bar{\alpha}\beta)^{n}\right).$
@@
where $\nu^{\prime}=\pi_{n-}^{\nu,\gamma_{\beta}^{\nu}}$ and $\gamma_{\beta}^{\nu^{\prime}}$ is an optimal discounted policy for $\nu^{\prime}$. As in [26], we can refer to these three components as transient cost, strategic measure cost, and approximation cost, respectively. The transient cost (17) is upper bound by
@@ 3.2. Approximations and Learning
Throughout the rest of this section, we assume that Assumption 1 holds. Earlier, we have already shown that if Assumption 1 is valid for the POMDP, then Assumption 9 is applicable for the belief-MDP. The following theorems assist in finding near-optimal policies for the average cost criterion; for further analysis see [29, Theorem 5] or [24, Theorem 7.3.6]. In particular, that a solution to the average cost optimality equation exists and this is arrived at via the vanishing discount method in our analysis earlier is critical for the applicability of the following result:
@@
Then, for every $\epsilon>0$, there exists $\beta_{\epsilon}\in(0,1)$ such that for $\beta\in[\beta_{\epsilon},1)$,
@@
This implies that the discounted cost optimal policy is near-optimal for the average cost criterion.
b) Under Assumption 9, let $\beta_{\epsilon}$ be chosen as above, such that with $h_{\beta}$ as given in (16),
@@
Since Assumption 1 for the POMDP implies Assumption 9 for the belief-MDP, as proven in the proof of Theorem 1.1, the above theorem also holds under Assumption 1. Therefore, under Assumption 1, a near-optimal policy for the discounted cost criteria is also near-optimal for the average cost criteria. The subsequent subsections will detail how to establish a near-optimal policy for discounted costs.
@@ 3.2.1. Near Optimality of Finite Models via Quantization for Average Cost
  • During exploration, every possible state-action pair in $\bar{Z}\times\mathbb{U}$ is visited an infinite number of times [^1] This can be relaxed and refined to concern those states in the support of an invariant measure under an exploration policy [31]..
@@ 3.2.2. Near Optimality of Finite Window Policies for Average Cost
(23) $z_{n-N}^{-}$ $=\operatorname{Pr}\left(X_{n-N}\in\cdot\mid y_{n-N-1},\ldots,y_{0},u_{n-N-1},\ldots,u_{0}\right),$
(24) $I_{n}^{N}$ =$=\left\{y_{n},\ldots,y_{n-N},u_{n-1},\ldots,u_{n-N}\right\},$
@@
Here, $\hat{\gamma}\in\hat{\Gamma}$, where $\hat{\Gamma}$ can be taken to be those Markov control policies under controlled states given with $\hat{z}_{n}$ defined in (22); that is, policies which map $(\hat{z}_{n},n)\to u_{n}$ for all $n\in\mathbb{Z}_{+}$.
With filter stability [32], as $N$ approaches infinity, the upper bound tends to zero. Given filter stability and a sufficiently large $N$, the policy $\tilde{\varphi}^{N}$ becomes near-optimal for the finite window belief MDP for discounted cost criteria. If $\beta$ is sufficiently large, this policy also becomes near-optimal for the average cost criterion.
@@
    Furthermore, we have that $P(Y_{t}=y|x)>0$ for every $x\in\mathbb{X}$, and thus during the exploration phase, every $(I,u)$ pair is visited infinitely often.
@@
[28], Theorem 4.1 provides sufficient conditions for uniform filter stability. If we have $\bar{\alpha}=(1-\tilde{\delta}(\mathcal{T}))(2-\delta(Q))<1$, then the filter is exponentially stable with coefficient $\bar{\alpha}$ for any control policy. Here, $\tilde{\delta}(\mathcal{T})=\inf_{u\in\mathcal{U}}\delta(T(\cdot|\cdot,u))$. See also an analysis via the Hilbert projective metric [33, 28].
@@ 4. Concluding Remarks
The average cost optimality of partially observable Markov decision process is a challenging problem. In this paper, we presented explicit and easily testable conditions for the existence of solutions to the average cost optimality equation where the state space is compact. A comparison with the related literature and several examples are presented. Notably, our paper appears to be the first one where a contraction result is presented and utilized for the optimality analysis. We finally presented several implications of our analysis and existence result on approximations, learning, and robustness.
@@
where $w_{y_{1}}=\left(z_{1}\left(z_{0}^{\prime},u,y_{1}\right)-z_{1}\left(z_{0},u,y_{1}\right)\right)$ which is a signed measure on $\mathbb{X}$. By the Measurable Selection Theorem 222[[34], Theorem 2][Kuratowski Ryll-Nardzewski Measurable Selection Theorem] Let $\mathbb{X},\mathbb{Y}$ be Polish spaces and $\Gamma=(x,\psi(x))$ where $\psi(x)\subset\mathbb{Y}$ be such that, $\psi(x)$ is closed for each $x\in\mathbb{X}$ and let $\Gamma$ be a Borel measurable set in $\mathbb{X}\times\mathbb{Y}$. Then, there exists at least one measurable function $f:\mathbb{X}\rightarrow\mathbb{Y}$ such that $\{(x,f(x)),x\in\mathbb{X}\}\subset\Gamma$. choose measurable
@@
Appendix B Proof of Lemma 2.2
We present a specialization of [24, Theorem 7.3.3] to the compact case as in the paper we have that $\mathcal{Z}$ is compact.
Let us first recall the Arzelà-Ascoli theorem.
Theorem B.1 ([35] Theorem 2.4.7).
Let $F$ be an equi-continuous family of functions on a compact space $\mathbb{X}$ and let $h_{n}$ be a sequence in $F$ such that the range of $f_{n}$ is compact. Then, there exists a subsequence $h_{n_{k}}$ which converges uniformly to a continuous function.
Proof of Lemma 2.2.
First, since the cost function $\tilde{c}$ is bounded (assume it is bounded by $M<\infty$) the expression $(1-\beta)J_{\beta}(z)$ is uniformly bounded by $M$ for every $\beta\in(0,1)$ and $z\in\mathcal{Z}$. By the Bolzano–Weierstrass theorem, for a fixed $z$ and any sequence $\beta\uparrow 1$, there exists a subsequence $\beta(k)\uparrow 1$ such that $\left(1-\beta(k)\right)J_{\beta(k)}\left(z\right)\rightarrow\rho^{*}$ for some $\rho^{*}$. Observe that for any $z\in\mathcal{Z}$
$\left(1-\beta(k)\right)J_{\beta(k)}(z)=\left(1-\beta(k)\right)\left(J_{\beta(k)}(z)-J_{\beta(k)}\left(z_{0}\right)\right)+\left(1-\beta(k)\right)J_{\beta(k)}\left(z_{0}\right),$
which, by the uniform boundedness of $h_{\beta(k)}(z)=J_{\beta(k)}(z)-J_{\beta(k)}\left(z_{0}\right)$, implies that the limit $\rho^{*}$ does not depend on $z$.
By Assumption 9-6, $h_{\beta(k)}$ is equicontinuous. By Theorem B.1, there exists a further subsequence of $h_{\beta(k)}$, $\left\{h_{\beta(k_{l})}\right\}$, which converges (uniformly on compact sets) to a continuous and bounded function $h$. Since $\mathbb{U}$ is compact and cost function is continuous and transition kernel is weakly continuous, we have by the Bellman equation:
$J_{\beta}(z)-J_{\beta}\left(z_{0}\right)$
(49) $=\min_{u\in\mathbb{U}}\left(c(z,u)+\beta\int\eta\left(dz^{\prime}\mid z,u\right)\left(J_{\beta}\left(z^{\prime}\right)-J_{\beta}\left(z_{0}\right)\right)-(1-\beta)J_{\beta}\left(z_{0}\right)\right)$
Take the limit in (B) along the subsequence $\beta(k_{l})$, we get
$h(z)$ $=\lim_{l}\min_{\mathbb{U}}\left[c(z,u)+\beta\left(k_{l}\right)\int_{\mathcal{Z}}h_{\beta(k_{l})}(y)\eta(dy\mid z,u)-(1-\beta(k_{l}))J_{\beta(k_{l})}\left(z_{0}\right)\right]$
$=\lim_{l}\min_{\mathbb{U}}\left[c(z,u)+\beta\left(k_{l}\right)\int_{\mathcal{Z}}h_{\beta(k_{l})}(y)\eta(dy\mid z,u)\right]-\rho^{*}$
From this we obtain
(50) $\lim_{l}\min_{\mathbb{U}}\left[c(z,u)+\beta\left(k_{l}\right)\int_{\mathcal{Z}}h_{\beta(k_{l})}(y)\eta(dy\mid z,u)\right]=h(z)+\rho^{*}$
We now show that in the above, the order of limit and minimization can be swapped: Using the compactness of $\mathbb{U}$, the continuity of
$c(z,u)+\beta(k_{l})\int_{\mathcal{Z}}h_{\beta(k_{l})}(y)\eta(dy|z,u)$
on $\mathbb{U}$, and the equicontinuity of $\left\{h_{\beta(k)}\right\}$ we can define a sequence $u_{l}$ such that:
$A_{l}:=c\left(z,u_{l}\right)+\beta\left(k_{l}\right)\int_{\mathcal{Z}}h_{\beta\left(k_{l}\right)}(y)\eta\left(dy\mid z,u_{l}\right)=$
$\min_{\mathbb{U}}\left[c(z,u)+\beta\left(k_{l}\right)\int_{\mathcal{Z}}h_{\beta\left(k_{l}\right)}(y)\eta(dy\mid z,u)\right]$
By the compactness of the action space $\mathbb{U}$, there exists a further subsequence such that $u_{l_{n}}\rightarrow u^{*}$ for some $u^{*}$ along this further subsequence. By weak continuity of the kernel, we then have that $\eta\left(dy\mid z,u_{l_{n}}\right)\rightarrow\eta(dy\mid z,u^{*})$. Since $h_{\beta\left(k_{l}\right)}$ is uniformly bounded, we have $\lim A_{l}=\lim A_{l_{n}}=c(z,u)+\int_{\mathcal{Z}}h(y)\eta(dy\mid z,u^{*})$.
On the other hand, for any fixed action $\bar{u}$, we know that
$c\left(z,\bar{u}\right)+\beta\left(k_{l}\right)\int_{\mathcal{Z}}h_{\beta\left(k_{l}\right)}(y)\eta\left(dy\mid z,\bar{u}\right)\geq A_{l}$
and taking the limit as $l\to\infty$, by the same argument,
$c(z,\bar{u})+\int_{\mathcal{Z}}h(y)\eta(dy\mid z,\bar{u})\geq\lim A_{l}=c(z,u)+\int_{\mathcal{Z}}h(y)\eta(dy\mid z,u).$
Therefore,
$\lim A_{l}=\min_{\mathbb{U}}\left[c(z,u)+\int_{\mathcal{Z}}h(y)\eta(dy\mid z,u)\right].$
Combining this with equation (50), we obtain
$h(z)+\rho^{*}=\min_{\mathbb{U}}\left[c(z,u)+\int_{\mathcal{Z}}h(y)\eta(dy\mid z,u)\right].$
Thus, we have found a bounded solution to the ACOE equation, completing the proof.